# Are We Truly Forgetting? A Critical Re-examination of Machine Unlearning Evaluation Protocols

### v1

@@ 3.2 Motivation: Why Do We Need Representation-Based Metrics?
Current machine unlearning algorithms predominantly rely on logit-based evaluation metrics to assess their effectiveness. These algorithms aim to achieve classification accuracy similar to models trained only on retain set ($\theta_{r}$), and based on these metrics, they are considered to have successfully accomplished unlearning. However, this evaluation raises a critical research question: Do existing unlearning evaluations adequately reflect successful unlearning from a representation perspective?
@@
t-SNE visualization. We first conducted an analysis on several unlearning baselines using t-SNE in Fig. 3, which provides a visualization of feature embeddings in a lower-dimensional space. While several unlearning algorithms (GA, RL, SalUn) achieve superior performance on logit-based evaluation metrics in small-scale settings, Fig. 3 reveals several important findings. First, (a) Original ($\theta_{o}$), (b) Retrain ($\theta_{r}$), and (c) Fine-Tuning (FT) models exhibit feature representations where each class forms distinct clusters. We expect that an effective unlearning algorithm should ideally produce feature representations similar to those of (b) ($\theta_{r}$). However, Figures (d)-(f) reveal that these approaches induce substantial collapse in feature representations, which deviate considerably from those of $\theta_{r}$, exposing a critical inconsistency. For t-SNE visualization results of other algorithms, see Fig. A in the Appendix.
CKA analysis. In Fig. 4, we compare the Centered Kernel Alignment (CKA) similarity values between the unlearned ($\theta_{u}$) and retrained ($\theta_{r}$) models (on the y-axis) against those between the unlearned and original ($\theta_{o}$) models (on the x-axis), using ImageNet-1k. Note that CKA quantifies the representational similarity between different models, providing a measure of how similarly two models “see” or interpret the given data. Intuitively, if two models are aligned in their feature representations, they can be considered to share a similar perspective on the data. In the context of successful unlearning, we expect the CKA similarity between the unlearned and retrained models to be higher than that between the unlearned and original models. However, our findings reveal a critical issue: all unlearning baselines show a higher similarity to the original model’s representations than to the retrained model’s representations. In fact, certain baselines, such as GA, SalUn, and RL, exhibit very low similarity to both the original and retrained models.
@@
Figure 4: CKA similarity analysis of various unlearning algorithms. The x-axis shows the similarity to the original model, and the y-axis represents the similarity to the retrained model. In an ideal scenario, algorithms should be positioned near the top-left corner, indicating high similarity to the retrained model and relatively low similarity to the original model. However, most algorithms exhibit a high similarity to the original model and limited alignment with the retrained model, suggesting that the transformation of representations during unlearning is suboptimal.
@@
Why does the unlearned model maintain high similarity to the original model? Several unlearning baselines show high logit-based metric scores (i.e., minimal deviation from the retrained model) in this experimental setup, while still maintaining a strong resemblance to the original model’s representations. We hypothesize that this occurs because unlearning algorithms tend to focus on modifying the final layer to optimize logit-based metrics, while leaving earlier representational layers largely unchanged. In Fig. 5, we compare the CKA similarities between unlearned models, where either all layers are tuned ($\theta_{u}$) or only the final layer is tuned ($\theta_{u}^{last}$), alongside the gap in logit-based AGL (Eq. 1) scores (y-axis). Ideally, the CKA similarity should be low (low x-values), which would correspond to a higher gap in AGL scores (high y-values). However, $\theta_{u}$ shows high similarity to $\theta_{u}^{last}$ while presenting a relatively small gap in AGL, demonstrating a crucial limitation in the current evaluation approach.
@@
Figure 5: The X-axis depicts the feature similarity between unlearned models with tuning all layers layer ($\theta_{u}$) versus only the final layer ($\theta_{u}^{last}$). Y-axis represents the gap in AGL scores between $\theta_{u}$ and $\theta_{u}^{last}$. The results indicate that $\theta_{u}$ modifies only the last layer while maintaining a similar representation space.
3.3 The Proposed Unified Evaluation Framework
To establish a rigorous and comprehensive evaluation of unlearning algorithms, we propose a unified framework that integrates both logit-based and representation-based metrics. This framework enables a more holistic assessment of unlearning performance by capturing both Feature Similarity and Feature Transferability—two fundamental aspects of an unlearned model’s behavior. First, Feature Similarity quantifies the alignment of feature representations between on the same dataset. To measure this, we employ Centered Kernel Alignment (CKA), which provides a robust similarity measure between the learned representations of both models. Second, Feature Transferability evaluates how well the unlearned model’s learned representations generalize to downstream tasks. We assess this by applying a $k$-Nearest Neighbors ($k$-NN) classifier to the encoder’s output features and assess performance on several downstream datasets. This allows us to measure how well features generalize beyond the original training distribution.
Unified metric.   To establish a comprehensive evaluation of unlearning algorithms, we propose a unified measure that integrates both logit-based and representation-based evaluations. For representation-based evaluation, we use diverse downstream datasets, $\mathcal{D}_{\text{down}}$, which differ from the original training distribution (e.g., ImageNet-1k). Inspired by transfer learning evaluation protocols [29], this approach enables us to assess both the quality and generalizability of learned representations. Specifically, we evaluate models on three downstream datasets: Office-Home [34], CUB [35], and DomainNet-126 [30]. Further details on these datasets can be found in Sec A.3 in the Appendix.
This metric evaluates the transfer learning ability of the given model’s encoder compared to that of the retrained model. In addition, we quantify the representational similarity between the given and retrained models for the given downstream dataset, by using $\mathrm{CKA}_{\text{AGR}}$:
$\mathrm{CKA}_{\text{AGR}}(\theta_{u},\mathcal{D}_{\text{down}}),$
which measures the representational similarity between the given and retrained models over the downstream dataset.
This formulation balances the impact of both feature divergence and transferability degradation, providing a more nuanced measure of unlearning effectiveness in a representation perspective.
$H\text{-}LR=\frac{2}{\frac{1}{AGL}+\frac{1}{AGR}}.$
This harmonic mean formulation ensures that both classification performance and representational divergence are jointly considered, making $H\text{-}LR$ a robust and comprehensive metric for evaluating unlearning algorithms.
Figure 6: AGL results for unlearning experiments on pretrained ResNet-50 models using CIFAR-10 and ImageNet-1k. We conduct unlearning on a pretrained model for one randomly selected class from CIFAR-10 and 100 randomly selected classes from ImageNet-1k. Notably, following the evaluation protocol of SalUn, we exclude test set accuracy when calculating AGL.
Figure 7: AGL and AGR results for unlearning experiments on pretrained ResNet-50 models using ImageNet-1k. We conduct unlearning on a pretrained model for 100 randomly selected classes from ImageNet-1k and report the corresponding AGL and AGR values. Notably, following the evaluation protocol of SalUn, we exclude test set accuracy when computing AGL.
@@ 3.4 Re-evaluation of Unlearning Baselines under Large-Scale Unlearning Scenarios
Small-scale vs. large-scale unlearning.   Several unlearning algorithms have demonstrated strong performance in small-scale settings, as evaluated by logit-based metrics. As shown in Fig. 6, both RL and SalUn achieve superior AGL on CIFAR-10. However, the figure also clearly highlights a significant performance degradation in ImageNet-1k, underscoring the challenges associated with scaling unlearning algorithms to larger datasets.
@@
Fig. 7 further presents both AGL and AGR scores in large-scale unlearning experiments. AGL scores, measured using ImageNet-1k, are accompanied by averaged AGR scores, which are calculated using multiple downstream datasets, $\mathcal{D}_{\text{down}}$. This figure emphasizes the limitations of both existing unlearning algorithms and the evaluation methods currently in use. First, three algorithms—GA, RL, and SalUn—demonstrate a significant drop in performance in the large-scale unlearning scenario. Second, all baseline algorithms consistently exhibit low AGR compared to AGL, indicating their failure to effectively unlearn from a representational perspective.
@@
Experimental results in large-scale unlearning.   Table 1 presents the experimental results for all baselines and evaluation methods in large-scale unlearning scenarios. Through logit-based evaluation and AGL scores, we observe that PL (which does not use $D_{r}$) achieves performance comparable to, or even surpassing, state-of-the-art algorithms that do utilize $D_{r}$ (i.e., DUCK, CU, SCAR, and SCRUB). This result marks a notable departure from the trends observed in small-scale unlearning experiments [11, 26, 2, 23]. In representation-based evaluation (measured by AGR), PL consistently outperforms other algorithms, achieving superior H-LR among the baselines. These results challenge prior expectations, showing that many state-of-the-art algorithms, such as SalUn, SCAR, and SCRUB, degrade in large-scale unlearning scenarios, performing worse from both logit and representation perspectives.
@@
Figure 8: Comparison of AGL and AGR across different unlearning baselines, using Random (R), Office-Home (OH), CUB (C), and DomainNet-126 (D) to select semantically similar 100 classes for top-class wise unlearning.
@@
Figure 9: AGL and AGR scores in Random and Top class-wise forgetting scenarios on PL. (a) AGL score archives consistently high in all settings. (b) AGR scores decrease, as the number of forgetting classes increases. This is more evident in top class-wise forgetting, where $\theta_{r}$ has a low similarity to $\theta_{o}$.
3.5 Experiments on Top Class-wise Forgetting
Why top class-wise forgetting is necessary?   In the random-class forgetting scenario, the CKA similarity between the retrained and original models reaches a high value of 88%, as shown in Fig. 4. We hypothesize that this high similarity arises from the presence of multiple alternative classes within ImageNet, which may preserve similar representational knowledge. Consequently, this results in a significant overlap between both the models, presenting challenges in verifying the effectiveness of unlearning.
To address this, we propose a novel top class-wise forgetting approach, where forgotten classes are selected based on semantic similarity criteria. This ensures that no alternative classes with similar representations remain in the retained data, leading to a substantial difference between $\theta_{r}$ and $\theta_{o}$. As shown in Fig. B in the Appendix, this approach minimizes CKA similarity, thereby facilitating a more reliable and straightforward evaluation of unlearning.
In selecting semantically similar classes, we leverage transfer learning techniques. Specifically, our top class-wise forgetting approach defines the forget set as the classes most similar to those in downstream datasets. We rank the classes and select the top 100 most similar ones for the forget set. For example, in the case of Office-Home, which consists of 65 classes, we identify the most similar classes to this dataset as the forget set. To complete the unlearning scenario, we extend the next 35 most similar classes, creating a 100-class forget set. If unlearning is successful, we expect to observe corresponding changes in performance on related classes in downstream tasks.
@@ 3.5 Experiments on Top Class-wise Forgetting
Need for considering both AGL and AGR, respectively.   Fig. 9(a) reveals that unlearning algorithms exhibit unexpected behaviors, which become evident when considering both AGR and AGL scores. Despite PL and DUCK often showing little disparity between these scores, there is an inherent need to assess these measures independently to fully understand the performance of the algorithms. To demonstrate this, we extended the evaluation of PL by performing experiments with 100 to 300 top class-wise unlearning classes. As the number of classes to forget increases, particularly in top class-wise forgetting, the nearest classes diverge significantly from the original ones. Fig. 9 illustrates that AGL scores remain largely unaffected by the number of forgetting classes. However, AGR scores notably decrease as the number of forgotten classes rises, indicating that PL retains substantial similarity to $\theta_{o}$ ’s representation while diverging from the retrained model. These findings not only highlight the discrepancy between the unlearning performed by these algorithms and the expected results but also emphasize the need for our proposed multi-faceted evaluation approach to better capture the effect of unlearning. Fig. C in the appendix shows the similar results on DUCK.
Note that we report detailed experimental and additional experimental results in Tables B-F in the Appendix.
4 Conclusion
This work introduces a new representation-based evaluation framework for large-scale unlearning, addressing the limitations of traditional logit-based metrics used in small-scale settings. Our experiments reveal that current unlearning algorithms struggle to scale in real-world scenarios, and logit-based metrics fail to capture the full impact of unlearning on model representations. Additionally, we show that random-class forgetting may not be enough to evaluate unlearning algorithms, while our top class-wise forgetting provides a more reliable and interpretable evaluation. These findings highlight the need for more robust and scalable unlearning methods, evaluated through our proposed framework.

### v2

@@ 3.2 Motivation: Why Do We Need Representation-Based Metrics?
Current machine unlearning algorithms predominantly rely on logit-based evaluation metrics to assess their effectiveness. These algorithms aim to achieve classification accuracy similar to model trained only on retain set ($\theta_{r}$), and based on these metrics, they are considered to have successfully accomplished unlearning. However, this evaluation raises a critical research question: Do existing unlearning evaluations adequately reflect successful unlearning from a representation perspective?
@@
t-SNE visualization.   We first conducted an analysis on several unlearning baselines using t-SNE in Fig. 3, which provides a visualization of feature embeddings from the encoder in a lower-dimensional space. While several unlearning algorithms (GA, RL, SalUn) achieve superior performance on logit-based evaluation metrics in small-scale settings, Fig. 3 reveals several important findings. First, (a) Original ($\theta_{o}$), (b) Retrain ($\theta_{r}$), and (c) Fine-Tuning (FT) models exhibit feature representations where each class forms distinct clusters. We expect that an effective unlearning algorithm should ideally produce feature representations similar to those of (b) ($\theta_{r}$). However, Figures (d)-(f) reveal that these approaches induce substantial collapse in feature representations, which deviate considerably from those of $\theta_{r}$. For all t-SNE visualization results, see Fig. 8 in the Appendix.
(a) CKA similarity of unlearning algorithms
(b) An analysis of $\theta_{u}^{last}$
@@
Figure 4: Two analyses using CKA similarity are presented: (a) CKA similarity analysis of various unlearning algorithms. The x-axis shows the similarity to $\theta_{o}$ and the y-axis represents the similarity to $\theta_{r}$. In an ideal scenario, algorithms should be positioned near $\theta_{r}$. However, most algorithms exhibit a high similarity to $\theta_{o}$, suggesting that the transformation of representations during unlearning is suboptimal. (b) The X-axis depicts the feature similarity between fully unlearned model ($\theta_{u}$) and only the final layer unlearned model ($\theta_{u}^{last}$). Y-axis represents the gap in AGL scores between $\theta_{u}$ and $\theta_{u}^{last}$. The results indicate that existing algorithms primarily modify only the last layer while maintaining the original representation space.
CKA analysis.   In Fig. 4(a), we compare the Centered Kernel Alignment (CKA) similarity values between the unlearned ($\theta_{u}$) and retrained ($\theta_{r}$) models (on the y-axis) against those between the unlearned and original ($\theta_{o}$) models (on the x-axis), using ImageNet-1K. Note that CKA quantifies the representational similarity between different models, providing a measure of how similarly two models “see” or interpret the given data. Intuitively, if two models are aligned in their feature representations, they can be considered to share a similar perspective on the data. In the context of successful unlearning, we expect the CKA similarity between the unlearned and retrained models to be higher than that between the unlearned and original models. However, our findings reveal a critical issue: all unlearning baselines show a higher similarity to the original model’s representations than to the retrained model’s representations. In fact, certain baselines, such as GA, SalUn, and RL, exhibit very low similarity to both the original and retrained models.
@@
Why does the unlearned model maintain high similarity to the original model?   Several unlearning baselines show high logit-based metric scores (i.e., minimal deviation from the retrained model) in our experimental setup, while still maintaining a strong resemblance to the original model’s representations. We hypothesize that this occurs because unlearning algorithms tend to focus on modifying the final layer to optimize logit-based metrics while leaving earlier representational layers largely unchanged. In Fig. 4(b), we compare the CKA similarities between unlearned models, where either all layers are tuned ($\theta_{u}$) or only the final layer is tuned ($\theta_{u}^{last}$), alongside the gap in logit-based AGL (Eq. 1) scores (y-axis). Ideally, the CKA similarity should be low (low x-values), which would correspond to a higher gap in AGL scores (high y-values). However, $\theta_{u}$ shows high CKA similarity to $\theta_{u}^{last}$ while presenting a relatively small gap in AGL, demonstrating a crucial limitation in the current evaluation approach.
@@
4 A Unified Benchmark Framework for Machine Unlearning
To establish a rigorous and comprehensive evaluation of unlearning algorithms, we propose a unified framework that integrates both logit-based and representation-based metrics. This framework enables a more holistic assessment of unlearning performance by capturing both Feature Similarity and Transferability—two fundamental aspects of an unlearned model’s behavior. First, Feature Similarity quantifies the alignment of feature representations of the original and unlearned models on the same dataset. To measure this, we employ Centered Kernel Alignment (CKA), which provides a robust similarity measure between the learned representations of both models. Second, Transferability evaluates how well the unlearned model’s learned representations generalize to downstream tasks. We assess this by applying a $k$-Nearest Neighbors ($k$-NN) classifier to the encoder’s output features and assess performance on several downstream datasets. This allows us to measure how well features generalize beyond the original training distribution.
@@ 4 A Unified Benchmark Framework for Machine Unlearning
Unified metric.   To establish a comprehensive evaluation of unlearning algorithms, we propose a unified measure that integrates both logit-based and representation-based evaluations. For representation-based evaluation, we use diverse downstream datasets, $\mathcal{D}_{\text{down}}$, which differ from the original train set distribution (e.g., ImageNet-1K). Inspired by transfer learning evaluation protocols [29], this approach enables us to assess both the quality and generalizability of learned representations. Specifically, we evaluate models on three downstream datasets: Office-Home [34], CUB [35], and DomainNet-126 [30]. More details of datasets are provided in Sec A.3 in the Appendix.
This metric evaluates the transfer learning ability of the given model’s encoder compared to that of the retrained model. In addition, we quantify the representational similarity between the given and retrained models for the given downstream dataset, by using $\mathrm{CKA}_{\text{AGR}}$: $\mathrm{CKA}_{\text{AGR}}(\theta_{u},\mathcal{D}_{\text{down}}),$ which measures the representational similarity between the given and retrained models over the downstream dataset.
This formulation balances the impact of both feature divergence and transferability degradation, providing a more nuanced measure of unlearning effectiveness from a representation perspective.
$H\text{-}LR=\frac{2}{\frac{1}{AGL}+\frac{1}{AGR}}.$
This harmonic mean formulation ensures that both classification performance and representational divergence are jointly considered, making $H\text{-}LR$ a robust and comprehensive metric for evaluating unlearning algorithms. Intuitively, a high H-LR score indicates that an unlearning algorithm achieves both output-level forgetting and meaningful representation-level divergence from the original model. In contrast, algorithms that modify only the classifier while leaving representations unchanged will score high on AGL but low on AGR, leading to a low H-LR. This unified metric penalizes such superficial unlearning.
@@ 4.1 Re-evaluation of Unlearning Baselines under Large-Scale Unlearning Scenarios
Small-scale vs. large-scale unlearning.   Several unlearning algorithms have demonstrated strong performance in small-scale settings, as evaluated by logit-based metrics. As shown in Fig. 5(a), both RL and SalUn achieve superior AGL on CIFAR-10. However, the figure also clearly highlights a significant performance degradation on ImageNet-1K, underscoring the challenges associated with scaling unlearning algorithms to larger datasets. Fig. 5(b) further presents both AGL and AGR scores in large-scale unlearning experiments. AGL scores, measured using ImageNet-1K, are accompanied by averaged AGR scores, which are calculated using multiple downstream datasets, $\mathcal{D}_{\text{down}}$. This figure emphasizes the limitations of both existing unlearning algorithms and the evaluation methods currently in use. First, three algorithms—GA, RL, and SalUn—demonstrate a significant drop in performance in the large-scale unlearning scenario. Second, all baseline algorithms consistently exhibit low AGR compared to AGL, indicating their failure to effectively unlearn from a representational perspective.
(a) Large vs. small scale
(b) AGL vs. AGR
Figure 5: (a) AGL results for unlearning experiments on pretrained ResNet-50 models using CIFAR-10 and ImageNet-1K. We conduct unlearning on a pretrained model for one randomly selected class from CIFAR-10 and 100 randomly selected classes from ImageNet-1K. (b) AGL and AGR results for unlearning experiments on ImageNet-1K pretrained ResNet-50 models. We conduct unlearning on a pretrained model for 100 randomly selected classes from ImageNet-1K and report the corresponding AGL and AGR values. Notably, following SalUn’s protocol, we excluded test set accuracy when calculating AGL in both (a) and (b)
@@
Experimental results in large-scale unlearning.   Table 1 presents the experimental results for all baselines and evaluation methods in our large-scale unlearning scenarios. Through logit-based evaluation and AGL scores, we observe that PL (which does not use $\mathcal{D}_{r}$) achieves performance comparable to or even surpassing state-of-the-art algorithms that do utilize $\mathcal{D}_{r}$ (i.e., DUCK, CU, SCAR, and SCRUB). This result marks a notable departure from the trends observed in small-scale unlearning experiments [11, 26, 2, 23]. In representation-based evaluation (measured by AGR), PL consistently outperforms other algorithms, achieving superior H-LR among the baselines. These results challenge prior expectations, showing that many state-of-the-art algorithms, such as SalUn, SCAR, and SCRUB, degrade in large-scale setup, performing worse from both logit and representation perspectives.
4.2 Experiments on Top Class-wise Forgetting
Why top class-wise forgetting is necessary?   In the random-class forgetting scenario, the CKA similarity between $\theta_{o}$ and $\theta_{r}$ reaches a high value of 88%, as shown in Fig. 4(a). We hypothesize that this high similarity arises from the presence of multiple alternative classes within ImageNet, which may preserve similar representational knowledge. Consequently, this results in a significant overlap between both models, presenting challenges in verifying the effectiveness of unlearning from a representational perspective.
To address this, we propose a novel top class-wise forgetting approach, where forgetting classes are selected based on semantic similarity criteria. This ensures that no alternative classes with similar representations remain in $\mathcal{D}_{r}$, leading to a substantial difference between $\theta_{r}$ and $\theta_{o}$. In other words, this setup makes unlearning more challenging, allowing for a more rigorous evaluation of unlearning algorithms. As shown in Fig. 9 in the Appendix, this approach minimizes CKA similarity between $\theta_{o}$ and $\theta_{r}$, thereby facilitating a more reliable and straightforward evaluation of unlearning.
Similarity-based Top-class Selection. In selecting semantically similar classes, we define $\mathcal{D}_{f}$ as the classes most similar to those in downstream datasets. To measure similarity, we extract feature maps for each class from the average pooling layer of ResNet-50 and compute cosine similarity between the class-wise averaged feature maps. We rank the classes and select the top 100 most similar ones for $\mathcal{D}_{f}$. For example, in the case of Office-Home, which consists of 65 classes, we identify the most similar ImageNet-1K classes to this dataset as the forget set. We extend the next 35 most similar classes, creating a 100-class forget set. If unlearning is successful, we expect to observe corresponding changes in performance on related classes in downstream tasks.
@@ 4.2 Experiments on Top Class-wise Forgetting
Figure 6: AGL and AGR comparison on unlearning baselines, using Random (R), Office-Home (OH), CUB (C), and DomainNet-126 (D) to select 100 similar classes for top class-wise unlearning.
@@
Figure 7: AGL and AGR scores in Random and Top class-wise forgetting scenarios on PL. (a) AGL score archives steadily high in all settings. (b) AGR scores decrease, as the number of forgetting classes increases. This is evident in top class-wise forgetting, where $\theta_{r}$ has a low similarity to $\theta_{o}$.
@@
Need for considering both AGL and AGR, respectively.   Fig. 6 reveals that unlearning algorithms exhibit unexpected behaviors, which become evident when considering both AGR and AGL scores. Despite PL and DUCK often showing little disparity between these scores, there is an inherent need to assess these measures independently to fully understand the performance of the unlearning algorithms. To demonstrate this, we extended the evaluation of PL by conducting experiments that unlearn 100 to 300 top semantically similar classes. As the number of forgetting classes increases, particularly in top class-wise forgetting, the nearest classes diverge significantly from the original ones. Fig. 7 illustrates that AGL scores remain largely unaffected by the number of forgetting classes. However, AGR scores notably decrease as the number of forgotten classes increases, indicating that PL still retains substantial similarity to $\theta_{o}$ ’s representation. These findings not only highlight the discrepancy between the unlearning performed by these algorithms and the expected results but also emphasize the need for our proposed multi-faceted evaluation approach. Fig. 10 in the appendix shows similar results on DUCK.
More experimental results are reported in the Appendix, including evaluations of alternative models (e.g., Swin-T [24]) under our experimental setup.
5 Concluding Remarks
Main contribution.   We propose a unified benchmark for evaluating machine unlearning under realistic, large-scale conditions. Our framework incorporates representation-based metrics (CKA, $k$-NN) and a novel Top Class-wise Forgetting scenario to reveal the limitations of current methods beyond traditional logit-based metrics. We show that many algorithms fail to significantly alter internal representations, despite showing high output-level scores. Our benchmark offers a scalable and reproducible evaluation protocol, laying the groundwork for more rigorous and reliable unlearning research. All code and datasets are publicly released to support community adoption.
Limitations.   Our benchmark is limited to classification tasks with accessible encoders and is not directly applicable to generative or black-box models. CKA and $k$-NN metrics, while informative, may not capture all semantic shifts. Future extensions should address dynamic unlearning and broader task types.
Broader impact.   Our benchmark enables more accountable and rigorous evaluation of unlearning algorithms, critical for privacy-sensitive applications such as healthcare or personalized AI. However, there’s a risk that future methods might overfit to benchmark metrics without achieving true forgetting. We advocate combining our benchmark with formal privacy guarantees and adversarial robustness checks to ensure safe deployment.